# Frostman random variables, entropy inequalities, and applications

### v4

@@ 5.2.2 Modifications to their argument
In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if we are only considering the situation $s\gg 1$.
@@ 7 Examples
In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Lebesgue measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
@@
Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have uniform distribution on $A$. In the following, we will prove several results. The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
@@
$\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X).$
In such situation, the number $l$ can be any non-negative integer, and $X$ is $(s,O(1))$-Frostman for any $s$ with $0\leq s\leq\frac{m}{m+l}$. So all the $s$ from $(0,1)$ are covered. Therefore, in the right-hand side of (2), the term $ns$ can not be replaced by $H_{n}(X)$.
@@
Calculation shows that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$. Indeed, the random variable $X$ is $(s,O(1))$-Frostman for any $0\leq s\leq\frac{m}{l+m}$. ∎
@@
It is sufficent to show that $H_{n}(X^{\prime 2})=(1+o(1))n$.
@@
Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in 7.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=\sqrt{X}$ and $\tilde{Y}=\sqrt{Y}$. By Example 7.1, one sees that
@@
Let $X$, $Y$ be the independent $(s,O(1))$-Frostman random variables defined in Example 7.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=1+X$, which is $(s,O(1))$-Frostman with range contained in $[1,2)$. Let $\tilde{Y}$ be an i.i.d. copy of $X$. Take $\phi(x,y)=(x+y)^{2}$, which is a degenerate quadratic form. By Example 7.1, one can deduce that
@@
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}=(1+o(1))ns.$
Hence, the requirement that $\phi$ is non-degenerate in Theorem 1.4 is essential.
@@
In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.4 if $(X,Y)$ is regarded as a marginally $(1/2,O(1);1/2,O(2^{m}))$-Frostman random variable, i.e., $s=1/2$. The reason is that the Frostman constant for $Y$ grows with $m$. More concretely, given $C$ of magnitude $2^{m}$, when $n=2m$ is of magnitude $\log C$, the two entropies may be small simultaneously. Therefore, in the conclusion of Theorem 1.4, it is essential to require $n>N$, where $N$ depends on the Frostman constant $C$. More concretely, the term $O_{s}(\log_{+}C_{1}C_{2})$ in the conclusion in Theorem 1.4’ is essential.
@@
This example is an variant of Example 7.4. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $G=\bigcup\nolimits_{j=0}^{2^{m}-1}\big[j2^{-m},\,j2^{-m}+2^{-2m}\big)^{2}$. Note that $X$ and $Y$ are dependent. Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$. We consider $m\leq n\leq 2m$.
Claim 1. The random variable $(X,Y)$ is marginally marginally $(1/2,O(1);1/2,O(1))$-Frostman, and jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(1))$-Frostman for any $(s_{1},s_{2})\neq(0,0)$.
@@
Now we consider the situation $n=2m$. By Claim 2, one has $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq n/2+O(1)$. In such a situation, the two entropies can be small simultaneously. Therefore, the requirement of independence of $X$ and $Y$ in Theorem 1.4 is essential.
@@
for some $\epsilon>0$. By Claim 2, the second and third entropies in above maximum-term nearly vanish. So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential. In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$. Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
In addition, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.38, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2. That is to say, the term $\log_{+}C$ in Theorem 2.38 is essential.
@@
$G^{\prime}=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
@@
Here $2^{-2m}\cdot G^{\prime}=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in G^{\prime}\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $G$.
@@
One can check that both $X$ and $Y$ have the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$. By Example 7.1, they are $(1/2,O(1))$-Frostman. ∎
@@
Note that $\mathcal{L}^{2}(A)=(2^{-4m})^{2}\cdot(2^{m})^{2}\cdot(2^{m})^{2}=2^{-4m}$. And $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be any Borel set, and $B(y,r)$ be any ball. Calculation reveals that
@@
For all the cases, it satisfies that

### v5

@@ 5.2.2 Modifications to their arguments
In the conclusions of Theorems 1.41.6, the term $C^{-O_{d,s}(1)}$ can be replaced by $C^{-O_{d}(1)}$ if we are only considering the situation $s\gg 1$.
7 Erdős–Szemerédi sum-product conjecture for adaptable sets
In this section, we translate Theorems 1.41.6 into purely discrete cardinality statements for finite sets. The key inputs are that uniformly $s$-adaptable sets automatically satisfy the required non-concentration hypothesis, whereas under the weaker assumption of $s$-adaptability one may pass to a large non-concentrated subset via a pruning argument.
7.1 Definitions and results
Definition 7.1.
Let $0<s<1$. Let $A\subseteq[0,1]$ be finite with $|A|=N$, and let $\delta:=N^{-1/s}$. Define the probability measure
$d\mu_{A}^{s}(x):=\frac{1}{N}\sum_{a\in A}\frac{1}{2\delta}\,\mathbf{1}_{[a-\delta,a+\delta]}(x)\,dx.$ (108)
We say that $A$ is $s$-adaptable with constant $C_{0}$ if
$I_{s}(\mu_{A}^{s}):=\iint_{\mathbb{R}^{2}}|x-y|^{-s}\,d\mu_{A}^{s}(x)\,d\mu_{A}^{s}(y)\leq C_{0},$ (109)
where $C_{0}$ is independent of $N$.
Definition 7.2.
We say that $A$ is uniformly $s$-adaptable with constant $C_{0}$ if for every $a\in A$,
$\sum_{\begin{subarray}{c}a^{\prime}\in A\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}\leq C_{0}\,N.$ (110)
We now list some examples of uniformly $s$-adaptable sets.
Let
$A_{N}=\left\{\frac{i}{N}:i=1,2,\dots,N\right\}\subseteq[0,1].$
Then, $A_{N}$ is $\delta$-separated because $1/N\geq\delta=N^{-1/s}$ for $s<1$. For any $a=i/N$,
$\sum_{a^{\prime}\neq a}|a-a^{\prime}|^{-s}\ll\sum_{k=1}^{N-1}\left(\frac{k}{N}\right)^{-s}=N^{s}\sum_{k=1}^{N-1}k^{-s}\leq N^{s}\left(1+\int_{1}^{N}x^{-s}\,dx\right)\ll_{s}N.$
So $A_{N}$ is uniformly $s$-adaptable with a constant $C_{0}=C_{0}(s)$. For non-lattice-structure sets, one can take
$A_{N}=\left\{\frac{i}{N}+\xi_{i}:i=1,\dots,N\right\}\cap[0,1],$
with $|\xi_{i}|\leq\frac{1}{100N}$ and $\xi_{i}$ monotone. Many other examples can be found in [14].
Theorem 7.3.
Let $d\geq 2$ and let $\rho_{1},\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$ and $C_{0}\geq 1$. Let $\epsilon=\epsilon(d,s)>0$ and $\delta_{0}=\delta_{0}(\rho_{1},\rho_{2},s,C_{0})>0$ be the constants from Theorem 1.4.
Let $A\subseteq[0,1]$ be any finite set satisfying $\delta:=(\#A)^{-1/s}\leq\delta_{0}$. Suppose that $A$ is $\delta$-separated and uniformly $s$-adaptable with constant $C_{0}$. If $G\subseteq A\times A$ satisfies the global density condition
$G\geq\delta^{\epsilon}(\#A)^{2},$ (111)
then
$(A+_{G}A)\;+\;\#(\rho_{1}(A)+_{G}\rho_{2}(A))\ \gg_{\rho_{1},\rho_{2},s}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}(\#A).$ (112)
Theorem 7.4.
Let $\phi$ be given as in (2), $0<s<1$ and $C_{0}\geq 1$. Let $\epsilon=\epsilon(d,s)>0$ or $\epsilon=\epsilon(d,\Xi,s)>0$, and $\delta_{0}=\delta_{0}(\phi,s,C_{0})>0$ be constants from Theorem 1.5.
Let $A\subseteq[0,1]$ be any finite set satisfying $\delta:=(\#A)^{-1/s}\leq\delta_{0}$. Suppose that $A$ is $\delta$-separated and uniformly $s$-adaptable with constant $C_{0}$. If $G\subseteq A\times A$ satisfies the two-sided minimum-degree conditions
$\{y\in A:(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\,\,(\forall x\in A),\quad\#\{x\in A:(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\,\,(\forall y\in A),$ (113)
then
$(A+_{G}A)+\#(\phi_{G}(A,A))\ \gg_{\phi,s}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,(\#A).$
Theorem 7.5.
Let $\phi$ be given as in (2), $0<s<1$, $C_{0}\geq 1$, and $c_{\mathrm{sep}}>1$. Let $\epsilon=\epsilon(d,s)>0$ or $\epsilon=\epsilon(d,\Xi,s)>0$ be the constant from Theorem 1.6. Let $A\subseteq[0,1]$ be any finite set and let $\delta:=(\#A)^{-1/s}$. Suppose that $A$ is $c_{\mathrm{sep}}\delta$-separated, and also $s$-adaptable with constant $C_{0}$.
Then, there exists a subset $A^{\sharp}\subseteq A$ with
$A^{\sharp}\gg_{s,c_{\mathrm{sep}}}\,\#A$
such that for every subset $B\subseteq A^{\sharp}$ with
$B\geq\delta^{\epsilon}\,(\#A),$
one has
$(B+B)+\#(\phi(B,B))\ \gg_{\phi,s,\mathrm{sep}}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,(\#A).$
The role of $A^{\sharp}$ is precisely to restore a non-concentration property not guaranteed by mere $s$-adaptability. Bounded discrete $s$-energy permits configurations of the form $A=A_{1}\cup A_{2}$, where $A_{1}$ is a large, highly concentrated block (e.g. a long arithmetic progression at spacing $\delta$) and $A_{2}$ is a dispersed remainder. Such sets can be $s$-adaptable yet fail the non-concentration estimate on intervals containing $A_{1}$. Taking $B=A_{1}$ then produces a large subset for which the hypotheses needed to force sum/polynomial expansion may fail, so one cannot expect a Theorem 7.5–type conclusion to hold for all $B\subseteq A$. Lemma 7.6 prunes away the points with large local energy and outputs a large subset $A^{\sharp}\subseteq A$ which is $\delta^{\sharp}$-separated and satisfies the required non-concentration condition at scale $\delta^{\sharp}$. This is why Theorem 7.5 is formulated for $B\subseteq A^{\sharp}$: once $A^{\sharp}$ satisfies non-concentration, one can apply the subset-stable expansion result (Theorem 1.6, with $A=A^{\sharp}$ and $\delta=\delta^{\sharp}$) to obtain the desired conclusion for every large $B\subseteq A^{\sharp}$.
7.2 Proof of Theorems 7.3 and 7.4
Proof of Theorem 7.3.
We verify the non-concentration hypothesis required for Theorem 1.4. Write N:=#AN:=\#A for simplicity. Let $I\subseteq\mathbb{R}$ be an interval with length $l\geq\delta$, and set $m:=\#(A\cap I)$. If $m\leq 1$ we are done. Otherwise, fix $a\in A\cap I$. Then for every $a^{\prime}\in A\cap I$ with $a^{\prime}\neq a$ we have $|a-a^{\prime}|\leq l$, hence
$C_{0}N\ \geq\ \sum_{\begin{subarray}{c}a^{\prime}\in A\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}\ \geq\ \sum_{\begin{subarray}{c}a^{\prime}\in A\cap I\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}\ \geq\ (m-1)\,l^{-s}.$
Therefore, $m\leq 1+C_{0}Nl^{s}$. Since $l\geq\delta$ implies $Nl^{s}\geq N\delta^{s}=1$, we obtain $m\leq 2C_{0}Nl^{s}$. So, the required non-concentration constant is $O(C_{0})$. Therefore, Theorem 1.4 gives
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\rho_{1}(A)+_{G}\rho_{2}(A))\ \gg_{\rho_{1},\rho_{2},s}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,\#A.$
Using $|S|\geq\mathcal{E}_{\delta}(S)$ for finite sets $S$, the theorem follows. ∎
Proof of Theorem 7.4.
As in the proof of Theorem 7.3, the non-concentration constant is $O(C_{0})$. Applying Theorem 1.5 gives
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\ \geq\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,\#A.$
Using $|S|\geq\mathcal{E}_{\delta}(S)$ for finite sets $S$, the theorem follows. ∎
7.3 Proof of Theorem 7.5
We first prove a pruning lemma that produces a large subset satisfying a non-concentration condition from $s$-adaptability.
Lemma 7.6.
Fix $0<s<1$, $C_{0}\geq 1$, and $c_{\mathrm{sep}}>1$. Let $A\subseteq[0,1]$ be any finite set with N:=#AN:=\#A, and let $\delta:=N^{-1/s}$. Suppose that $A$ is $c_{\mathrm{sep}}\,\delta$-separated, and is $s$-adaptable with constant $C_{0}$. Let $K\geq\max\left\{\frac{1}{1-c_{sep}^{-s}},2\right\}$, and define
$N_{K}:=\Big\lceil\Big(1-\frac{1}{K}\Big)N\Big\rceil,\qquad\delta_{K}:=N_{K}^{-1/s}.$
Then, there exists a subset $A_{K}\subseteq A$ with $A_{K}=N_{K}$ such that
(i) $A_{K}$ is $\delta_{K}$-separated with $\delta_{K}\leq c_{\mathrm{sep}}\delta$,
(ii) for every interval $I\subseteq\mathbb{R}$ of length $l\geq\delta_{K}$,
$(A_{K}\cap I)\leq C_{\mathrm{nc}}\,l^{s}\,\#A_{K},$ (114)
where $C_{\mathrm{nc}}\leq C(s)\,K\,C_{0}$ for some constant $C(s)>0$ depending only on $s$.
Proof.
For $a\in A$, define the discrete energy
$E(a):=\sum_{\begin{subarray}{c}a^{\prime}\in A\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}.$
We first compare the continuous energy $I_{s}(\mu_{A}^{s})$ with the discrete sum. Since $A$ is $c_{\mathrm{sep}}\delta$-separated, for any distinct $a,a^{\prime}\in A$ and any $x\in[a-\delta/4,a+\delta/4]$, $y\in[a^{\prime}-\delta/4,a^{\prime}+\delta/4]$ we have
$|x-y|\leq|a-a^{\prime}|+\frac{\delta}{2}\leq\frac{3}{2}\,|a-a^{\prime}|.$
Using the definition (108), we have
$\iint|x-y|^{-s}\,d\mu_{A}^{s}(x)\,d\mu_{A}^{s}(y)\;\geq\;\Big(\frac{1}{2\delta N}\Big)^{2}\Big(\frac{\delta}{2}\Big)^{2}\Big(\frac{3}{2}\Big)^{-s}\,|a-a^{\prime}|^{-s}=c(s)\,\frac{|a-a^{\prime}|^{-s}}{N^{2}},$
where $c(s):=2^{-4}(3/2)^{-s}>0$. Summing over all ordered pairs $a\neq a^{\prime}$ gives
$I_{s}(\mu_{A}^{s})\geq c(s)\,\frac{1}{N^{2}}\sum_{\begin{subarray}{c}a,a^{\prime}\in A\\ a\neq a^{\prime}\end{subarray}}|a-a^{\prime}|^{-s}=c(s)\,\frac{1}{N^{2}}\sum_{a\in A}E(a).$
Since $I_{s}(\mu_{A}^{s})\leq C_{0}$, we conclude that
$\frac{1}{N}\sum_{a\in A}E(a)\leq C^{\prime}(s)\,C_{0}\,N,$
for $C^{\prime}(s):=c(s)^{-1}$.
A direct computation shows that at most $N/K$ points satisfy $E(a)>K\,C^{\prime}(s)\,C_{0}\,N$. Hence there is a subset $A^{\prime}\subseteq A$ with $A^{\prime}\geq(1-1/K)N$ such that
$E(a)\leq K\,C^{\prime}(s)\,C_{0}\,N\qquad\text{for all }a\in A^{\prime}.$
Choose any subset $A_{K}\subseteq A^{\prime}$ with $A_{K}=N_{K}$. Then, $A_{K}$ is $\delta_{K}$-separated, since $A_{K}\subseteq A$ and $\min_{a\neq a^{\prime}}|a-a^{\prime}|\geq c_{\mathrm{sep}}\delta\geq\delta_{K}$.
To prove (114), we use similar arguments as in the proof of Theorem 7.3 to get, for any interval $I$ with length $l\geq\delta_{K}$,
$(A_{K}\cap I)\leq 1+K\,C^{\prime}(s)\,C_{0}\,N\,l^{s}\leq C(s)\,K\,C_{0}\,l^{s}\,N_{K}$
for some $C(s)>0$. This completes the proof. ∎
Proof of Theorem 7.5.
Let $\epsilon>0$ be the exponent from Theorem 1.6. Since Theorem 1.6 remains valid if $\epsilon$ is replaced by a smaller positive number, we may assume that
$0<\epsilon<s.$ (115)
Write N:=#AN:=\#A. Choose
$K_{*}\ :=\ \left\lceil\max\left\{2,\ \frac{1}{1-c_{\mathrm{sep}}^{-s}}\right\}\right\rceil\qquad\text{and set}\qquad N_{*}\ :=\ \left\lceil\left(1-\frac{1}{K_{*}}\right)N\right\rceil.$
Apply Lemma 7.6 with $K=K_{*}$, and let
$A^{\sharp}\ :=\ A_{K_{*}}\subseteq A,\qquad\#A^{\sharp}=N_{*},\qquad\delta^{\sharp}\ :=\ (N_{*})^{-1/s}.$
Then, Lemma 7.6 gives:
(i) $A^{\sharp}$ is $\delta^{\sharp}$-separated and $\delta^{\sharp}\leq c_{\mathrm{sep}}\delta$;
(ii) for every interval $I\subseteq{\mathbb{R}}$ with length $l\geq\delta^{\sharp}$,
$(A^{\sharp}\cap I)\ \leq\ C_{\mathrm{nc}}\,l^{s}\,\#A^{\sharp},\qquad\text{with}\qquad C_{\mathrm{nc}}\ \leq\ C(s)\,K_{*}\,C_{0}.$ (116)
Moreover,
$A^{\sharp}=N_{*}\ \geq\ \left(1-\frac{1}{K_{*}}\right)N.$ (117)
Let $B\subseteq A^{\sharp}$ with $B\geq\delta^{\epsilon}\#A$. We claim that $B$ is large enough to apply Theorem 1.6 at scale $\delta^{\sharp}$. Indeed, using (115) and $A^{\sharp}\leq\#A$,
$(\delta^{\sharp})^{\epsilon}\#A^{\sharp}=(N_{*})^{1-\epsilon/s}\leq N^{1-\epsilon/s}=\delta^{\epsilon}(\#A)\leq\#B.$
Hence
[…]
We may therefore apply Theorem 1.6 to the $\delta^{\sharp}$-separated set $A^{\sharp}$, using the non-concentration estimate (116), and the subset $B$ satisfying (118). This implies
$\mathcal{E}_{\delta^{\sharp}}(B+B)\ +\ \mathcal{E}_{\delta^{\sharp}}(\phi(B,B))\ \gg_{\phi,s}\ C_{\mathrm{nc}}^{-O_{s}(1)}\,(\delta^{\sharp})^{-\epsilon}\,\#A^{\sharp}.$
Since $S\geq\mathcal{E}_{\delta^{\sharp}}(S)$ for any finite set $S$, we obtain
$(B+B)\ +\ \#(\phi(B,B))\ \gg_{\phi,s}\ C_{\mathrm{nc}}^{-O_{s}(1)}\,(\delta^{\sharp})^{-\epsilon}\,\#A^{\sharp}.$ (119)
It remains to compare $(\delta^{\sharp})^{-\epsilon}\#A^{\sharp}$ with $\delta^{-\epsilon}\#A$. By (117),
$(\delta^{\sharp})^{-\epsilon}(\#A^{\sharp})=(N_{*})^{1+\epsilon/s}\geq\left(1-\frac{1}{K_{*}}\right)^{1+\epsilon/s}N^{1+\epsilon/s}=\left(1-\frac{1}{K_{*}}\right)^{1+\epsilon/s}\,\delta^{-\epsilon}\,(\#A).$
Also, from (116) we have $C_{\mathrm{nc}}\leq C(s)K_{*}C_{0}$, so
$C_{\mathrm{nc}}^{-O_{s}(1)}\ \geq\ C_{0}^{-O_{s}(1)}\,c_{0}(s,c_{\mathrm{sep}})$
for some $c_{0}(s,c_{\mathrm{sep}})>0$ depending only on $s$ and $c_{\mathrm{sep}}$. Combining with (119) gives
$(B+B)\ +\ \#(\phi(B,B))\ \geq_{\phi,s}\ C_{0}^{-O_{s}(1)}\,\Big(c_{0}(s,c_{\mathrm{sep}})\left(1-\tfrac{1}{K_{*}}\right)^{1+\epsilon/s}\Big)\,\delta^{-\epsilon}\,(\#A).$
Finally, write
$c(s,c_{\mathrm{sep}})\ :=\ \min\left\{1-\frac{1}{K_{*}},\ \ c_{0}(s,c_{\mathrm{sep}})\left(1-\frac{1}{K_{*}}\right)^{1+\epsilon/s}\right\}.$
Then, (117) implies $A^{\sharp}\geq c(s,c_{\mathrm{sep}})\,(\#A)$, and the last display implies the desired lower bound. This completes the proof. ∎
@@ 8 Examples
In this section, we will show details of some examples. A short summary is listed in advance in Table 1 below.
Example Construction What it demonstrates
8.1 $X,Y$ i.i.d., uniform on $A=\bigcup\nolimits_{j=0}^{2^{m}-1}[j2^{-m},j2^{-m}+2^{-m-l})$ $\bullet$ lower bounds in Lem 2.32(i) are tight $\bullet$ Thm 1.3 fails when $s=0$ or 11 $\bullet$ nsns cannot be replaced by $H_{n}(X)$ in Thms 1.2 and 1.3 $\bullet$ single term $H_{n}(X+Y)$ can be small
8.2 $\tilde{X}=\sqrt{X}$, $\tilde{Y}=\sqrt{Y}$ from above $\bullet$ single term $)H_{n}(\phi(\tilde{X},\tilde{Y}$)) can be small
8.3 $\tilde{X}=1+X$, $\tilde{Y}=Y$ from above $\phi(x,y)=(x+y)^{2}$ $\bullet$ Thm 1.3 fails if $\phi$ can be reduced to a univariate function
8.4 $X$ uniform on intervals $Y$ uniform on $0,2^{-2m}] \bullet$ Thm 1.3 fails when $C$ grows with $n$
8.5 $(X,Y)$ uniform on $\bigcup_{j=0}^{2^{m}-1}[j2^{-m},j2^{-m}+2^{-2m}]^{2}$ $\bullet$ Thm 1.3” fails without independence $\bullet$ $\frac{1}{2}(H_{n}(X)+H_{n}(Y))$ in Thm 1.2 cannot be discarded $\bullet$ $\log^{+}C$ is essential in Thm 2.38 $\bullet$ jointly $(1/4;1/4)$ but marginally $(1/2;1/2)$
8.6 $(X,Y)$ uniform on carefully constructed $G$ $\bullet$ conditionally (1/4; 1/4) but jointly (1/2; 1/2)
8.7 $X$, $Y$ as in 8.1 $\bullet$ $H_{n}(X+\sqrt{2}Y)-H_{n}(X)$ are much larger than $H_{n}(X+Y)-H_{n}(X)$ $\bullet$ Thm 2.23 fails when $k$ and $l$ are $\mathbb{Q}$-independent.
Table 1: Summary of examples demonstrating sharpness and necessity of conditions in main theorems. Examples 7.1–7.5 show various conditions are essential; Examples 7.5 and 7.6 demonstrate the Frostman hierarchy structure; Example 7.7 shows rational arithmetic is crucial.
Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Lebesgue measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
@@
Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that $X$ and $Y$ are independent, and both have uniform distribution on $A$. In the following, we will prove several results. The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
@@
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\right\}$
$\qquad=\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X).$
In such situations, the number $l$ can be any non-negative integer, and $X$ is $(s,O(1))$-Frostman for any $s$ with $0\leq s\leq\frac{m}{m+l}$. So all the $s$ from $(0,1)$ are covered. Therefore, in the right-hand side of (1) and (3), the term $ns$ can not be replaced by $H_{n}(X)$.
@@
Computation shows that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$. Indeed, the random variable $X$ is $(s,O(1))$-Frostman for any $0\leq s\leq\frac{m}{l+m}$. ∎
@@
It is sufficient to show that $H_{n}(X^{\prime 2})=(1+o(1))n$.
@@
Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in 8.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=\sqrt{X}$ and $\tilde{Y}=\sqrt{Y}$. By Example 8.1, one sees that
@@
Let $X$, $Y$ be the independent $(s,O(1))$-Frostman random variables defined in Example 8.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=1+X$, which is $(s,O(1))$-Frostman with range contained in $[1,2)$. And $\tilde{X}$, $Y$ are independent. Take $\phi(x,y)=(x+y)^{2}$, which is a degenerate quadratic form. By Example 8.1, one can deduce that
@@
$\max\{H_{n}(\tilde{X}+Y),\,H_{n}(\phi(\tilde{X},Y))\}=(1+o(1))ns.$
Therefore, in Theorem 1.3, it is essential to require that $\phi$ cannot be reduced to a univariate function.
@@
In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.3 if $(X,Y)$ is regarded as independent marginally $(1/2,O(1);1/2,O(2^{m}))$-Frostman random variable, i.e., $s=1/2$. The reason is that the Frostman constant for $Y$ grows with $m$. More concretely, given $C$ of magnitude $2^{m}$, when $n=2m$ is of magnitude $\log C$, the two entropies may be small simultaneously. Therefore, in the conclusion of Theorem 1.3, it is essential to require $n>N$, where $N$ depends on the Frostman constant $C$. In other words, the term $O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})$ in the conclusion in Theorems 1.3’ and 1.3” is essential.
@@
This example is a variant of Example 8.4. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $G=\bigcup\nolimits_{j=0}^{2^{m}-1}\big[j2^{-m},\,j2^{-m}+2^{-2m}\big)^{2}$. Note that $X$ and $Y$ are dependent. Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$. We consider $m\leq n\leq 2m$.
Claim 1. The random variable $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman, and jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(1))$-Frostman for any $(s_{1},s_{2})\neq(0,0)$.
@@
Now we consider the situation $n=2m$. By Claim 2, one has $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq n/2+O(1)$. In such a situation, the two entropies can be small simultaneously. Therefore, the requirement of independence of $X$ and $Y$ in Theorem 1.3” is essential.
@@
for some $\epsilon>0$. By Claim 2, the second and third entropies in above maximum term nearly vanish. So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential. In this example, it is $m+O(1)$, which is larger than $ns=m/2$.
In addition, for the random variable $W$, its Frostman constant $C$ can grow with $m$. Consider $n=2m$ in Theorem 2.38. If $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2. That is to say, the term $\log_{+}C$ in Theorem 2.38 is essential.
@@
$G^{\prime}=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},j2^{-m}+(i+1)2^{-2m}).$
@@
Here $2^{-2m}\cdot G^{\prime}=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in G^{\prime}\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $G$. It is not hard to see that $X$ and $Y$ are dependent.
@@
One can check that both $X$ and $Y$ have the uniform distribution on the set $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-2m}\right)$. By Example 8.1, they are $(1/2,O(1))$-Frostman. ∎
@@
Note that $\mathcal{L}^{2}(G)=(2^{-4m})^{2}\cdot(2^{m})^{2}\cdot(2^{m})^{2}=2^{-4m}$. And $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be any Borel set, and $B(y,r)$ be any ball. Calculation reveals that
@@
For all the cases, we have that